# Renormalization group improved cosmology in the presence of a stiff matter era

### v1

@@ Abstract
In Phys. Rev. D 92 (2015) 103004, simple analytical solutions of the Friedman equations were obtained for a universe having stiff matter component in the early universe together with a dark matter, and a dark energy component. In this analysis, the universe is considered to be made of a dark fluid which behaves as a stiff matter in the early phase of the universe (when the internal energy dominates). It is also more logical to consider quantum gravitational effects in the early phase of the cosmological evolution. In this analysis, following Phys. Rev. D 65 (2002) 043508, we consider renormalization group improved modified Friedmann equations where the Newton’s gravitational constant ($G$) and the cosmological constant ($\Lambda$) flows with the momentum scale $k$ of the universe. In Phys. Rev. D 65 (2002) 043508, $k$ is identified with either $\frac{1}{t}$ or with the inverse of the scale factor of the universe $\left(\frac{1}{a(t)}\right)$. In this work, we obtain the momentum scale in terms of the scale factor of the universe when a stiff-matter era is present in the early time regime ($t$ being smaller than the Planck time). It is observed that for a universe undergoing a stiff matter era, radiation era, and matter era, inflation is absent in the early time regime of the universe when the flow of the Newton’s gravitational constant and cosmological constant is under consideration. Using the identification of the momentum scale with the scale factor of the universe, we then explore the era $t>t_{\text{Pl}}$ which indicates a primarily matter dominated era with accelerated expansion due to the presence of dark energy. Finally, considering the total equation of state as a combination of linear equation of state along with a polytropic equation of state, we observe that after the Planck-time the universe can undergo an inflationary phase and we find out that the inflation is enhanced by quantum gravitational effects arising due to the consideration of renormalization group approach to quantum gravity.
@@ I Introduction
The initial notion of the universe was that it is homogeneous, isotropic and static at the same time. Einstein in [1, 2], used his general relativistic field equations and applied it to cosmology. He took into account the existence of homogeneity and isotropy in the universe. He has considered a static universe which was the standard notion upto then where the universe had no beginning and it exists forever. In order to obtain a static solution he needed to consider a cosmological constant in the field equations. In the same year de Sitter also proposed a static cosmological solution with the astrophysical bodies and system considered to be test particles, in capable of affecting the geometry of the spacetime [3]. In his paper, he predicted that the light emitted from the sources will be redshifted because of the repulsive nature generated due to the cosmological constant and with increasing distance time will slow down which is known as the de Sitter’s effect. The static model of Einstein could not explain redshift whereas de-Sitter’s model consisted of no matter leading to certain inconsistencies. First Friedmann [4, 5] and then Lanczos [6], Weyl [7], and Lemaître [8, 9] proposed non-stationary solutions of the universe. This particular model of the universe explained redshift without resorting to any adhoc assumptions. Robertson also gave a reinterpretation of the de Sitter solution by taking dynamics into consideration [10, 11]. In [12], Eddington argued and proved that the static model of the universe proposed by Einstein is unstable which was also supported by the observations made by Edwin Hubble [13]. As a result it was evident that the universe is homogeneous, isotropic but non-static in nature. Finally, in [14, 15], Lemaître proposed that the universe must have started from a singular point (a singularity) as it is expanding. This paved the way for the Big Bang model of the universe. One problem with the Big Bang model is that at $t=0$ the energy density goes to infinity. This is also known as the singularity problem of the universe. Such problems only occur because at such high energy scale, quantum gravity effects should come into effect rather than general classical field equations. The Big Bang theory has several problems like the flatness problem, horizon problem, and monopole problem which can be solved using a inflation in the evolution of the cosmos [26, 27, 28]. One of the ways to deal with the horizon problem and the flatness problem is by introducing renormalized group improved cosmology. Renormalization group improvement in black hole spacetime was first done in [20, 21]. In [22], renormalization group technique was implemented to determine the leading order quantum gravity corrections in the well known Friedmann-Lemaître-Robertson-Walker (FLRW) cosmology. The basic step is to consider an effective average action $\Gamma_{k}[g_{\mu\nu}]$ where $k$ is a momentum scale with $\Gamma_{k}$ perfectly taking care of all tree level gravitational phenomenon as well as all loop corrections (for momentum of the order $k$). If $\Gamma_{k}$ is treated as a function of the momentum scale $k$, then $\Gamma_{k}$ implies a renormalization group trajectory which is placed in the space of all action functionals. One can determine such a trajectory by solving the “flow equations”. Non-perturbative approximate solutions for the renormalization group equations can be obtained via the method of “truncation” where from the infinite dimensional space of all action functionals, one projects the renormalization group flow onto some finite dimensional relevant subspace. This truncation method converts the functional renormalization group equations to a system of ordinary differential equations where the ordinary differential equations involves a finite set of generalized coupling constants. For a renormalization group improved FLRW cosmological model, the generalized coupling constants are the Newton’s gravitational constant $G(k)$ and cosmological constant $\Lambda(k)$ where the coupling constants depend on the momentum scale $k$. In [23], the differential equations governing the dynamics of $G(k)$ and $\Lambda(k)$ depending on the momentum scale $k$ were derived and the solutions were further discussed in [24, 21]. If one takes a very high value of $k$ ($k\rightarrow\infty$), then the dimensionless Newton’s gravitational constant $g(k)=k^{2}G(k)$ approaches a $UV$ attractive non-Gaussian fixed point $g_{*}^{UV}$ which also implies that gravity is asymptotically free ($G(k)\rightarrow 0$ for $k\rightarrow\infty$). The Weinberg fixed point in $2+\epsilon$-dimensions is similar to the current fixed point for 3+1-dimensional quantum gravity theory [25]. The renormalization group improvement is implemented by replacing $G$ by $G(k)$ and $\Lambda$ by $\Lambda(k)$ in the Einstein’s field equations. In [22] it was argued that the renormalization group improved solutions are “natural” implying the absence of the need of any fine tuning and the solutions are free of particle horizons for a broad class of equations of state. Therefore, these solutions get rid of the flatness problem and horizon problem without the need of an inflation era. There also has been several literature based on renormalized group improved cosmology in recent analysis [26, 27, 28].
@@
It is important to note all of the analyses above has considered the existence of a radiation era, a dark matter era. The current accelerated expansion of the universe is governed by a dark energy dominated era, also known as the late inflation. The cosmological model first proposed by Zel’dovich [29], brought in the possibility of a primordial stiff matter era. In this model, the primordial universe is assumed to be constructed of a very cold gas of baryons governed by an equation of state $P=\varepsilon$. He showed that for such a stiff matter era the speed of sound is almost equal to the speed of light [30]. As the energy density corresponding to the stiff matter era $\varepsilon_{\text{s}}$ is proportional to $a(t)^{-6}$, it will dominate before the radiation, dark matter, and dark energy era. After the success of the hot big bang model the model proposed by Zel’dovich was abandoned although, as argued in [31], a primordial stiff matter era is of great physical interest. A stiff-matter era implicates the initial state of the universe was in a state of inactivity rather than in a state of chaos [31]. For relativistic scalar field theory, if the kinetic term dominates the potential term of the relativistic scalar field, the scalar field starts to behave as a stiff fluid. In [32], a fully relativistic treatment of Bose-Einstein condensate dark matter/ stiff fluid was conducted which explicitly showed the existence of a primordial stiff matter era, followed by a radiation era and a matter era where the radiation era existed because of a self-interacting stiff fluid. Using a hydrodynamical representation of the Klein-Gordon-Einstein equation the same model proposed in [32] was analyzed in [33]. In the same year in [34], cosmology with a general stiff matter era was considered where for a stiff matter with negative energy density was shown to result in bouncing solutions of the universe as observed in loop quantum cosmology [35].
@@
Our paper is organized as follows. In section(II), we discuss briefly the classical Friedmann equations and flow improved Friedmann equations. Then in section(III), we discuss the the modified equation of state using the relativistic thermodynamical approach. Next, in section(IV) and section(V), we obtain the solutions of the flow improved Friedmann equations for different combinations of the various components present in the universe in the fixed point regime and in the beyond fixed point regime or the perturbative regime. In section(VI), we discuss inflation for a polytropic equation of state. Finally, in section(VII), we summarize our results.
II Background of the analysis
II.1 Friedmann equations
We begin our study by considering that the universe is isotropic and homogeneous. The isotropic and homogeneous model of the universe is described by the Friedmann-Lemaître-Robertson-Walker metric or the FLRW metric and the corresponding line element reads [^1] The speed of light is taken to be unity which has been later restored.
@@ II.1 Friedmann equations
with $u_{\mu}=\{1,0,0,0\}$ in a co-moving frame of reference. The Einstein’s field equations with non-vanishing cosmological constant reads
$\begin{split}R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R=8\pi GT_{\mu\nu}-\Lambda g_{\mu\nu}~{}.\end{split}$ (3)
Setting $\mu=0$ and $\nu=0$ in eq.(3), we arrive at the first Friedmann equation which is given by
$\left(\frac{\dot{a}}{a}\right)^{2}=\frac{8\pi G\varepsilon}{3}+\frac{\Lambda}{3}-\frac{3\kappa}{a^{2}}~{}.$ (4)
Similarly using the $\{r,r\}$ component from eq.(3) and making use of eq.(4), we obtain the second Friedmann equation as [^2] The $\{\theta,\theta\}$ and $\{\phi,\phi\}$ components give the same analytical form as eq.(5).
$\frac{\ddot{a}}{a}=-\frac{4\pi G}{3}(\varepsilon+3P)+\frac{\Lambda}{3}+\frac{\kappa}{a^{2}}~{}.$ (5)
Again $\nabla^{\mu}G_{\mu\nu}=0$, which along with the metric compatibility condition gives from eq.(3)
$\nabla^{\mu}T_{\mu\nu}=0~{}.$ (6)
Using eq.(2) in the above equation, we obtain
@@
As experimental evidence shows that the universe is almost flat, we can proceed with $\kappa=0$, and write down the modified Friedmann equations as
$\begin{split}H^{2}&=\frac{8\pi G\varepsilon}{3}+\frac{\Lambda}{3}\\ \frac{\ddot{a}}{a}&=-\frac{4\pi G}{3}(\varepsilon+3P)+\frac{\Lambda}{3}\\ \frac{d\varepsilon}{dt}&=-\frac{3\dot{a}}{a}(\varepsilon+P)\end{split}$ (8)
where the Hubble constant is defined by $H\equiv\frac{\dot{a}}{a}$.
II.2 Renormalization group improved cosmology
In this subsection, we shall briefly discuss the renormalization group approach to quantum gravity and cosmology developed in [21, 22]. The primary way to such an approach is to consider an effective average action $\Gamma_{k}[g_{\mu\nu}]$ which depends on the flow of the momentum scale $k$. As a result of this renormalization group improved approach, the Newtonian gravitational constant $G$ and the cosmological constant $\Lambda$, flows with the momentum scale $k$ and can be written respectively as $G(k)$ and $\Lambda(k)$. In [21, 23, 24], the differential equations governing the dynamics of $G(k)$ and $\Lambda(k)$ were obtained. As the momentum scale $k$ approaches infinity, $G(k)$ approaches zero which implies that gravity is asymptotically free for such a theory. As $k\rightarrow\infty$, the dimensionless Newtonian constant $g(k)$, where $g(k)=k^{2}G(k)$, approaches an UV-attractive fixed point $g_{*}^{UV}$. As a result one can express the flowing gravitational constant for very high values of $k$ as
$G(k)=\frac{g_{*}^{UV}}{k^{2}}~{}.$ (9)
The cosmological constant $\Lambda(k)$ for very high values of $k$ with an UV-attractive fixed point $\lambda_{*}^{UV}$ reads
$\Lambda(k)=\lambda_{*}^{UV}k^{2}~{}.$ (10)
Here, $G(k)$ and $\Lambda(k)$ follow the asymptotic solutions given in eq.(s)(9,10), if and only if $k$ is larger than the Planck mass [22]. The renormalization group improved Friedmann equations read
$\frac{\ddot{a}}{a}$ $=-\frac{4\pi G(k)}{3}(\varepsilon+3P)+\frac{\Lambda(k)}{3}~{}.$ (12)
Doing a time derivative of eq.(11) and making use of eq.(12), we obtain
From the covariant conservation of stress energy tensor $\nabla^{\mu}T_{\mu\nu}=0$ condition, we still get back eq.(7), in presence of the flow of $G$ and $\Lambda$ with the momentum scale $k$. For a homogeneous and isotropic universe the scale factor $k$ can be a function of the cosmological time only [22]. Comparing eq.(7) with eq.(13), we then obtain a constraint condition involving $\Lambda$, $G$, and $\varepsilon$ as
@@ II.2 Renormalization group improved cosmology
If $k\rightarrow 0$ (for $t\gg t_{\text{Pl}}$), the dynamical relations of $G$ and $\Lambda$ follows the following equations [22]
@@
with the threshold function $\Phi_{n}^{m}(w)$ defined as [22] ($m=1,2,\ldots$)
@@
with $R^{(0)}$ denoting the cut-off function. In eq.(s)(16,17), $G_{0}$ and $\Lambda_{0}$ denotes the value of the Newton’s gravitational constant and cosmological constant in the present time. Using eq.(14), and eq.(s)(16,17), one can obtain the form of the energy density in terms of $k$ as (upto lowest order contribution in $k$) [36]
@@
Our primary aim in this work is to make use of the renormalization group improved Friedmann equations to investigate the implications of the existence of a stiff-matter era in the early universe and analyse its underlying implications.

### v3

@@ Abstract
In Phys. Rev. D 92 (2015) 103004, simple analytical solutions of the Friedman equations were obtained for a universe having stiff matter component in the early universe together with a dark matter, and a dark energy component. In this analysis, the universe is considered to be made of a dark fluid which behaves as a stiff matter in the early phase of the universe (when the internal energy dominates). It is also more logical to consider quantum gravitational effects in the early phase of the cosmological evolution. In this analysis, following Phys. Rev. D 65 (2002) 043508, we consider renormalization group improved modified Friedmann equations where the Newton’s gravitational constant ($G$) and the cosmological constant ($\Lambda$) flows with the momentum scale $k$ of the universe. In Phys. Rev. D 65 (2002) 043508, $k$ is identified with either $\nicefrac{{1}}{{t}}$ or with the inverse of the scale factor of the universe $\left(\nicefrac{{1}}{{a(t)}}\right)$. In this work, we obtain the momentum scale in terms of the scale factor of the universe when a stiff-matter era is present in the early time regime ($t$ being smaller than the Planck time). It is observed that for a universe undergoing a stiff matter era, radiation era, and matter era, inflation is absent in the early time regime of the universe when the flow of the Newton’s gravitational constant and cosmological constant is under consideration. Using the identification of the momentum scale with the scale factor of the universe, we then explore the era $t>t_{\text{Pl}}$ which indicates a primarily matter dominated era with accelerated expansion due to the presence of dark energy. Finally, considering the total equation of state as a combination of linear equation of state along with a polytropic equation of state, we observe that after the Planck-time the universe can undergo an inflationary phase and we find out that the inflation is enhanced by quantum gravitational effects coming due to consideration of renormalization group improvement in the Newton’s gravitational constant as well as the cosmological constant. Finally, we propose a new energy density that shows the transition between a primordial stiff matter era followed by an inflation era just after the Planck time, and then the transition from inflation to radiation dominated reheating phase around a second transition point in this renormalization group improved cosmological model.
Stiff matter, Renormalization group improved cosmology, Inflation
@@ I Introduction
The initial notion of the universe was that it is homogeneous, isotropic and static at the same time. Einstein in AEinstein1 ; AEinsteinReview , used his general relativistic field equations and applied it to cosmology. He took into account the existence of homogeneity and isotropy in the universe. He has considered a static universe which was the standard notion upto then where the universe had no beginning and it exists forever. In order to obtain a static solution he needed to consider a cosmological constant in the field equations. In the same year de Sitter also proposed a static cosmological solution with the astrophysical bodies and system considered to be test particles, in capable of affecting the geometry of the spacetime deSitter1 . In his paper, he predicted that the light emitted from the sources will be redshifted because of the repulsive nature generated due to the cosmological constant and with increasing distance time will slow down which is known as the de Sitter’s effect. The static model of Einstein could not explain redshift whereas de-Sitter’s model consisted of no matter leading to certain inconsistencies. First Friedmann Friedmann1 ; Friedmann2 and then Lanczos Lanczos , Weyl Weyl , and Lemaître Lemaitre ; Lemaitre2 proposed non-stationary solutions of the universe. This particular model of the universe explained red-shift without resorting to any adhoc assumptions. Robertson also gave a reinterpretation of the de Sitter solution by taking dynamics into consideration Robertson1 ; Robertson2 . In Eddington , Eddington argued and proved that the static model of the universe proposed by Einstein is unstable which was also supported by the observations made by Edwin Hubble Hubble . As a result it was evident that the universe is homogeneous, isotropic but non-static in nature. Finally, in Lemaitre3 ; Lemaitre4 , Lemaître proposed that the universe must have started from a singular point (a singularity) as it is expanding. This paved the way for the Big Bang model of the universe. One problem with the Big Bang model is that at $t=0$ the energy density goes to infinity. This is also known as the singularity problem of the universe. Such problems only occur because at such high energy scale, quantum gravity effects should come into effect rather than general classical field equations. The Big Bang theory has several problems like the flatness problem, horizon problem, and monopole problem which can be solved using a inflation in the evolution of the cosmos RGCosmo0 ; RGCosmo1 ; RGCosmo2  [^1] One can also look into non-singular bouncing cosmological models for a physical analysis of spacetime defects Battista .. One of the ways to deal with the horizon problem and the flatness problem is by introducing renormalized group improved cosmology. Renormalization group improvement in black hole spacetime was first done in BonanoReuter99 ; BonanoReuter . In BonanoReuter2 , renormalization group technique was implemented to determine the leading order quantum gravity corrections in the well known Friedmann-Lemaître-Robertson-Walker (FLRW) cosmology. The basic step is to consider an effective average action $\Gamma_{k}[g_{\mu\nu}]$ where $k$ is a momentum scale with $\Gamma_{k}$ perfectly taking care of all tree level gravitational phenomenon as well as all loop corrections (for momentum of the order $k$). If $\Gamma_{k}$ is treated as a function of the momentum scale $k$, then $\Gamma_{k}$ implies a renormalization group trajectory which is placed in the space of all action functionals. One can determine such a trajectory by solving the “flow equations”. Non-perturbative approximate solutions for the renormalization group equations can be obtained via the method of “truncation” where from the infinite dimensional space of all action functionals, one projects the renormalization group flow onto some finite dimensional relevant subspace. This truncation method converts the functional renormalization group equations to a system of ordinary differential equations where the ordinary differential equations involves a finite set of generalized coupling constants. For a renormalization group improved FLRW cosmological model, the generalized coupling constants are the Newton’s gravitational constant $G(k)$ and cosmological constant $\Lambda(k)$ where the coupling constants depend on the momentum scale $k$. In Reuter , the differential equations governing the dynamics of $G(k)$ and $\Lambda(k)$ depending on the momentum scale $k$ were derived and the solutions were further discussed in Souma ; BonanoReuter . If one takes a very high value of $k$ ($k\rightarrow\infty$), then the dimensionless Newton’s gravitational constant $g(k)=k^{2}G(k)$ approaches a $UV$ attractive non-Gaussian fixed point $g_{*}^{UV}$ which also implies that gravity is asymptotically free ($G(k)\rightarrow 0$ for $k\rightarrow\infty$). The Reuter fixed point in $2+\epsilon$-dimensions is similar to the current fixed point for 3+1-dimensional quantum gravity theory Weinberg0 . The renormalization group improvement is implemented by replacing $G$ by $G(k)$ and $\Lambda$ by $\Lambda(k)$ in the Einstein’s field equations. In BonanoReuter2 it was argued that the renormalization group improved solutions are “natural” implying the absence of the need of any fine tuning and the solutions are free of particle horizons for a broad class of equations of state. Therefore, these solutions get rid of the flatness problem and horizon problem without the need of an inflation era. There also has been several literature based on renormalized group improved cosmology in recent times RGCosmo0 ; RGCosmo1 ; RGCosmo2 .
Recently in Donoghue , it was argued that the asymptotic safety approach discussed above has some fundamental flaws. It was argued in Donoghue that the asymptotic safety approch runs into some conflicts with exact analytical results in low energies. It has been argued that just by applying renormalization group improvement to Newton’s coupling the one-loop correction to gravity mediated scattering amplitude cannot be obtained and one needs to improve the “standard practice” of renormalization group improvement by implementing a Lorentzian signature instead of using the Euclidean signature. As has been discussed in ReplytoDonoghue , the implementation of a metric with Euclidean signature has the benefit of easily defining the direction of the Renormalization group flow. For a quantum field theory in flat spacetime one calculate the propagators in ther Euclidean signature and then restore the Lorentzian model by implementing a Wick rotation of the Euclidean time. In curved spacetime however, the analytical continuation from the Euclidean to Lorentzian signature is not straighforward and rather has some very severe obstacles. One of them being the complex structure of the graviton propagator which will restrict one to use the analytical continuation from Euclidean to Lorentzian signature. The momentum regulation in this process may break the global spacetime symmetries or result in generating new poles AnalyticalContinuationIssue . Hence the extension of analytical continuation in asymptotic safety gravity has not yet been implemented. Hence, we stick with the standard practice of this functional renormalization-group approach and will use the standard techniques thoroughly used in several literatures.
@@
It is important to note all of the analyses above has considered the existence of a radiation era, a dark matter era. The current accelerated expansion of the universe is governed by a dark energy dominated era, also known as the late inflation. The cosmological model first proposed by Zel’dovich Zeldovich , brought in the possibility of a primordial stiff matter era. In this model, the primordial universe is assumed to be constructed of a very cold gas of baryons governed by an equation of state $P=\varepsilon$. He showed that for such a stiff matter era the speed of sound is almost equal to the speed of light Zeldovich2 . As the energy density corresponding to the stiff matter era $\varepsilon_{\text{s}}$ is proportional to $a(t)^{-6}$, it will dominate before the radiation, dark matter, and dark energy era. After the success of the hot big bang model the model proposed by Zel’dovich was abandoned although, as argued in NatureStiff , a primordial stiff matter era is of great physical interest. A stiff-matter era implicates the initial state of the universe was in a state of inactivity rather than in a state of chaos NatureStiff . For relativistic scalar field theory, if the kinetic term dominates the potential term of the relativistic scalar field, the scalar field starts to behave as a stiff fluid. In BECSF , a fully relativistic treatment of Bose-Einstein condensate dark matter/ stiff fluid was conducted which explicitly showed the existence of a primordial stiff matter era, followed by a radiation era and a matter era where the radiation era existed because of a self-interacting stiff fluid. Using a hydrodynamical representation of the Klein-Gordon-Einstein equation the same model proposed in BECSF was analyzed in ChavanisSuarez . In the same year in Chavanis0 , cosmology with a general stiff matter era was considered where for a stiff matter with negative energy density was shown to result in bouncing solutions of the universe as observed in loop quantum cosmology AshtekarSingh .
@@
Our paper is organized as follows. In section(II), we discuss briefly the classical Friedmann equations and flow improved Friedmann equations. Then in section(III), we discuss the the modified equation of state using the relativistic thermodynamical approach. Next, in section(IV) and section(V), we obtain the solutions of the flow improved Friedmann equations for different combinations of the various components present in the universe in the $k>m_{Pl}$ regime and in the $k<m_{Pl}$ regime or the perturbative regime. In section(VI), we have investigated whether it is possible to get bouncing solutions of the universe. Next, in section(VII), we discuss inflation for a polytropic equation of state. Finally, in section(VIII), we summarize our results.
II Brief review of the background model
II.1 Renormalization group improved cosmology and the Friedmann equations
We start our discussion by considering the Friedmann-Lemaître-Robertson-Walker metric or the FLRW metric for an isotropic and homogeneous model of the universe which is given by [^2] The speed of light is taken to be unity.
@@ II.1 Renormalization group improved cosmology and the Friedmann equations
with $u_{\mu}=\{1,0,0,0\}$ in a co-moving frame of reference. As experimental evidence shows, our universe is almost flat. Hence, without loss of any generality, we can proceed with $\kappa=0$. As our main focus is to investigate a stiff-matter dominated cosmology where renormalization group flows plays an important role, we shall start by briefly discussing the renormalization group approach to quantum gravity and cosmology developed in BonanoReuter ; BonanoReuter2 . The primary way to such an approach is to consider an effective average action $\Gamma_{k}[g_{\mu\nu}]$ which depends on the flow of the momentum scale $k$. As a result of this renormalization group improved approach, the Newtonian gravitational constant $G$ and the cosmological constant $\Lambda$, flows with the momentum scale $k$ and can be written respectively as $G(k)$ and $\Lambda(k)$. In BonanoReuter ; Reuter ; Souma , the differential equations governing the dynamics of $G(k)$ and $\Lambda(k)$ were obtained. As the momentum scale $k$ approaches infinity, $G(k)$ approaches zero which implies that gravity is asymptotically free for such a theory. As $k\rightarrow\infty$, the dimensionless Newtonian constant $g(k)$, where $g(k)=k^{2}G(k)$, approaches an UV-attractive fixed point $g_{*}^{UV}$. As a result one can express the flowing gravitational constant for very high values of $k$ as Reuter
$G(k)=\frac{g_{*}^{UV}}{k^{2}}~{}.$ (3)
The cosmological constant $\Lambda(k)$ for very high values of $k$ with an UV-attractive fixed point $\lambda_{*}^{UV}$ reads Reuter
$\Lambda(k)=\lambda_{*}^{UV}k^{2}~{}.$ (4)
Here, $G(k)$ and $\Lambda(k)$ follow the asymptotic solutions given in eq.(s)(3,4), if and only if $k$ is larger than the Planck mass BonanoReuter2 . The renormalization group improved Friedmann equations read BonanoReuter2 (for $\kappa=0$)
$\frac{\ddot{a}}{a}$ $=-\frac{4\pi G(k)}{3}(\varepsilon+3P)+\frac{\Lambda(k)}{3}$ (6)
where $k\equiv k(t)$ is a function of $t$ as will be identified in the subsequenct discussion. Doing a time derivative of eq.(5) and making use of eq.(6), we obtain
Implementing the covariant conservation of stress energy tensor which is given by the condition $\nabla^{\mu}T_{\mu\nu}=0$, we get back the standard continuity equation given as
@@
In the presence of the flow of $G$ and $\Lambda$ with the momentum scale $k$. For a homogeneous and isotropic universe the scale factor $k$ can be a function of the cosmological time only BonanoReuter2 . Comparing eq.(7) with the standard continuity equation in Friedmann-Lemaître-Robertson-Walker cosmology, we obtain a constraint condition involving $\Lambda$, $G$, and $\varepsilon$ as RGCosmo2
@@
This indeed is a very important relation in a sense that one can express the energy-density of the universe in terms of the momentum scale introducing a flow-like condition of the energy density itself. It is imortant to note that if $G(k)$ and $\Lambda(k)$ are considered as time-dependent constants then eq.(7) reduces to the standard continuity equation in eq.(8) turning eq.(9) into a “0=0” identity. If $k\rightarrow 0$ (for $t\gg t_{\text{Pl}}$), the dynamical relations of $G$ and $\Lambda$ follows the following equations BonanoReuter2
@@
with the threshold function $\Phi_{n}^{m}(w)$ defined as BonanoReuter2 ($m=1,2,\ldots$)
@@
with $R^{(0)}$ denoting the cut-off function. In eq.(s)(11,12), $G_{0}$ and $\Lambda_{0}$ denotes the value of the Newton’s gravitational constant and cosmological constant in the present time. Using eq.(9), and eq.(s)(11,12), one can obtain the form of the energy density in terms of $k$ as (upto lowest order contribution in $k$) Cosmologyfinal
@@
Our primary aim in this work is to make use of the renormalization group improved Friedmann equations to investigate the implications of the existence of a stiff-matter era in the early universe and analyse its underlying implications. [^3] This section is purely intended for a brief pedagogical review of the background model consisting of results from previous works that we shall use for our current analysis.